Circular No. 6065

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994Y IN NGC 5371 A. Clocchiatti, R. Garcia-Lopez, E. S. Barker, W. Wren, and J. C.

Wheeler, University of Texas, report: “A fully-reduced spectrogram (range 380–640 nm, resolution 1.3 nm) of SN 1994Y, obtained on Aug. 27.15 UT at the 2.7-m telescope of McDonald Observatory, shows the lines of the Balmer series and Na D in emission superimposed on a blue continuum. In the low-S/N spectrum, the emission lines do not show a P-Cyg absorption counterpart with the possible exception of Na D, but consist of a narrow core and a broad base, both centered at the velocity of NGC 5371. The FWHM of the broader components implies an expansion velocity of ∼ 2600 km/s, while the FWZI implies a maximum expansion of ∼ 7000 km/s. SN 1994Y apparently belongs to the ‘Seyfert 1’ subclass of type-II supernovae (Filippenko 1989, A.J. 97, 726), called type-IIn by Schlegel (1990, MNRAS 244, 269). Continuous monitoring of this event is strongly recommended.”

X. J. Jiang, W. Liu, and J. Y. Hu, Beijing Observatory, communicate: “A spectrogram (range 450–700 nm, resolution 0.53 nm/pixel) of SN 1994Y has been obtained using the 2.2-m telescope at Beijing Observatory, Xinglong, on Aug. 25. The spectrum shows strong Balmer emission, confirming it to be a type-II supernova. The wavelengths of the Hα and Hβ emission lines coincide with the radial velocity of the host galaxy NGC 5371 (2550 km/s). The emission lines are narrow compared with normal type-II supernovae; the FWHM of Hα and Hβ are ∼ 1200 km/s. Another dominating wide emission feature peaks at 592.9 nm, possibly due to He I (587.6 nm) and Na I (589.3 nm). A feature at 673.0 nm is from He I (667.8 nm). The spectrum is similar to that of SN 1994W, but without narrow absorption features at the blue side of the Balmer emission lines.”

A. Boattini and M. Tombelli, Florence, provide an accurate position for this object: α = 13^{h}53^{m}30^{s}.57, δ = +40^{o}42′32′′.8 (equinox 1950.0).

LT DELPHINI R. Passuello and S. Saccavino, Associazione Friulana di Astronomia e

Meteorologia (AFAM), Udine; and U. Munari, Asiago Astrophysical Observatory, report their photometry of this symbiotic star, obtained with the AFAM 0.45-m telescope on Aug. 26.9 UT: V = 12.22, B − V = +0.56, U − B = −0.74 (± 0.05 mag in all colors). Comparison with archival data collected on 1990 June 23, when the star was at V = 13.13, B −V = +1.26, U − B = −0.48 (Munari et al. 1992, A.Ap. Suppl. 93, 383), suggests that this star is currently undergoing an unexpected bright phase.

1994 August 29 Daniel W. E. Green